OpenStudy (anonymous):

Which pairs are solutions to the equation? 5xy+9=44 A.(1,7) and (7,1) B.(7,1) and (2,3) C.(1,7) and (4,11) D.(4,11) and (2,3)

OpenStudy (nincompoop):

plug in one of the coordinates and see if the equation is true

OpenStudy (anonymous):

5(1)(7)+9=44 5(7)(1)+9=44 5(7)+9=44 35(1)+9=44 35+9=44 35+9=44 44=44 44=44 It would be a

OpenStudy (anonymous):

I got it dont worry Thanks any way guys

OpenStudy (nincompoop):

ahuh...

OpenStudy (nincompoop):

lol